4x+7y=29
5x+8y=34
Elimination method

Answer by Boreal(15235)

4x+7y=29
5x+8y=34; multiply top by -5 and bottom by 4 to eliminate the x
-20x-35y=-145
20x+32y=136
-3y=-9, adding
y=3
substitute into the first for x
4x+7(3)=29
4x=8
x=2
(2,3)
Check in the second
5(2)+8(3)=34 checks.